OpenStudy (parissfrance09):

Line A has a slope of 2 and passes through the point (3, 15). Line B has a slope of 2 and passes through the point (4, 17). Which of the following statements is true about the system of equations represented by line A and line B? A.The system of equations has no real solutions. B.The system of equations has an infinite number of real solutions. C.The system of equations has exactly one real solution. D.The system of equations has exactly two real solutions.

OpenStudy (parissfrance09):

@pooja195 @I_Ask_Alot @Hero @SamsungFanBoy @Lexi_Loves @Callisto @Michele_Laino @Beast-Mode

OpenStudy (aravindg):

First can you write the line equations?

OpenStudy (parissfrance09):

how?

OpenStudy (parissfrance09):

@AravindG

OpenStudy (michele_laino):

for example, in order to write the equation of both lines, we can apply the subsequent equation: \(y-y_0=m(x-x_0)\)

OpenStudy (parissfrance09):

I havent learned that yet

OpenStudy (aravindg):

If a line has slope m and passes through point (a,b) Then its line equation is y-b=m(x-a)

OpenStudy (aravindg):

Never too late to learn!

OpenStudy (parissfrance09):

okay so what do i do next

OpenStudy (michele_laino):

hint: since the lines above have the same slope, then such lines are parallel. What can you conclude?

OpenStudy (parissfrance09):

i dont know im reallly confused

OpenStudy (michele_laino):

The general theory says that parallel lines, have the same slope

OpenStudy (michele_laino):

a solution of the system is the intersection point, if it exists, of the two lines above. Now parallel lines, according to the geometry of Euclid, have not any intersection point